OpenStudy (anonymous):

on interval 0>x>2pi sec(x)=-1

OpenStudy (anonymous):

\[secx = -1 => \frac{1}{cosx} = -1 \]

OpenStudy (anonymous):

Are you able to do it now? :)

OpenStudy (anonymous):

honestly, im completely clueless all together. haha

OpenStudy (anonymous):

well do you know that \[secx = \frac{1}{cosx}\]?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

then where are you haveing problems?

OpenStudy (anonymous):

what about the PI

OpenStudy (anonymous):

I'm just not too sure what exactly the question wants me to do honestly

OpenStudy (anonymous):

well looks like the question wants you to solve for \(x\) why not try and look in your texbook?

OpenStudy (anonymous):

We dont have one yet, since this is summer work

OpenStudy (anonymous):

well i know pi is =3.1415926535897932384626433832795028841971693993751058209749445923078164062862089986280348253421170679 but it is 3.14 for shot

OpenStudy (anonymous):

that doesnt solve much :P

OpenStudy (anonymous):

ya i know just trying to help

OpenStudy (anonymous):

0>x>2pi sec(x)=-1 this is hard what grad is this

OpenStudy (anonymous):

AP calculus

OpenStudy (anonymous):

no no; not hard at all :) \[secx = -1 => \frac{1}{cosx} = -1 => 1 = -cosx => -cosx = 1 => cosx = -1 => x = \pi \] this is only trig not calculus yet :)

OpenStudy (anonymous):

ah must be review problems then. Is that the full answer?

OpenStudy (anonymous):

lol the answer got cut off. why not you try and solve it ? :)

OpenStudy (anonymous):

not funy

OpenStudy (anonymous):

I have no idea what im doing and this is the last problem i have thats due tomorrow lol

OpenStudy (anonymous):

:(

OpenStudy (anonymous):

need this answer :\